### v1

@@
Abstract
@@
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
*Equal contribution
@@ 1 Introduction
Today, large-language model (LLM) inference has outpaced training in growth. LLM inference powers a wide range of applications, from interactive customer support and code generation to retrieval-based document analysis and agentic workflows. However, the LLM inference systems have not caught up—latency (both response latency and generation speed) and cost are now the bottleneck.
A key limitation of today’s LLM inference systems is that each user query is independently processed by one instance of the inference engine, without any data being reused across queries or inference engines. This means that a LLM query’s lifecycle, including computation and I/O operations, happens in the GPUs and the GPU memory of one inference engine. Each query takes a token sequence as input and another token sequence as output, and once a query completes, all outputs or intermediate states generated during inference are discarded.
This design, albeit simple, leads to redundant computation and resource underutilization, and two notable industry proposals attempted to address them (Figure 1).
Cross-query caching to avoid redundant compute:  As each query is treated in isolation, the LLM must prefill the same context (prefix) each time the context is used, and the repeated prefill compute is expensive and causes slow response, as users cannot see the first token before the prefill completes. Yet, such redundant computation could be avoided if we persist KV cache of the prefix—which LLM created internally after it prefills the prefix—beyond the lifecycle of a query and reuse the KV cache when subsequent queries reuse the prefix (e.g., the same prompt template or retrieved document chunk). By avoiding redundant computation, cross-query context (prefix) caching significantly lowers both time-to-first-token (TTFT) and overall GPU resource consumption during the prefill phase.
Prefill–decode disaggregation for higher utilization:  Co-locating the prefill phase (compute-bound, throughput-oriented) and the decode phase (memory-bound, latency-sensitive) in the same inference engine causes resource underutilization as the GPUs must be overprovisioned to ensure a consistently low decoding latency without being interrupted by prefill. In response, the industry attempted to decouple the throughput-oriented prefill processing from latency-sensitive decoding Such prefill-decode (PD) diaggregation architecture batches all queries’ prefill phase on one set of inference nodes and transfers their KV caches to another set of inference nodes that run the decoding phase only, thus mitigating tail decoding latency under high concurrency.
@@
To make these ideas practical, the LLM inference systems must be augmented with new KV cache semantics. In particular, inference engines should support the new interface that extracts KV caches from a normal inference call and re-loads KV caches into subsequent queries on demand. The system also must allow the extracted KV caches to be stored persistently and transferred across distributed inference engines. Most importantly, for such interface extensions to be practical, KV cache extraction, re-loading, storage, and transfer must be efficient, and the new interfaces must remain compatible with rapidly evolving inference engines such as vLLM Kwon et al. (2023b) and SGLang Zheng et al. (2024).
@@
The remainder of this paper details the motivation (§2), LMCache architecture and key design choices (§3–§6), experimental evaluation (§8), and deployment experiences (§9).
2 Motivation
2.1 Inference Cost and Delay as the Bottleneck
The past two years have witnessed explosive growth in the deployment of LLMs across diverse enterprise applications. As the user base and query volume of LLM-based applications increase, the cost and delay of inference, rather than training, emerges as the primary bottleneck to scaling LLM serving to production scale.
Inference cost:  Besides the simple fact that LLMs are queried by more users, the increasing input length and output length also contribute to raising inference cost, as each input token and each output token incur an additional cost to process. Four trends are behind the increase in input and output lengths. (i) More interactions by each user with LLMs build a longer user history, which will be prepended to future LLM inputs as the context. (ii) LLMs increasingly take multimodal inputs, such as images and videos, which typically will be transformed to a long sequence of tokens before being fed to LLMs. (iii) Newer LLMs often have longer context windows, allowing context/prompt engineers to simply stuff more tokens into LLM inputs. (iv) Finally, reasoning models produce unusually long output, which might be used as input to subsequent queries in one workflow.
Even with steady improvement on GPU price and performance, serving a trillion-token-scale workload can incur millions of dollars in annual compute expenditure Nie et al. (2024); Databricks Research (2024). Redundant computation across queries (e.g., recomputing the same prompt prefix hundreds of times) exacerbates these costs Zhang et al. (2024); Caylent (2024).
Inference delay:  In interactive applications, tail latency (e.g., 95th/99th percentile) is as critical as median latency. Slow responses in the worst case, often due to resource contention or large context processing, directly degrade user experience and can lead to user abandonment, in much the same way video streaming stalls affect viewer experience Liu et al. (2024a).
In particular, empirical observations from industry reports and academic studies indicate that:
  • $\bullet$
    Time-to-first-token (TTFT) can be dominated by the prefill phase for long-context inputs, especially when context lengths exceed 8–16k tokens.
  • $\bullet$
    Inter-token latency (ITL) remains low for single-stream decoders, but rises sharply when GPU compute is shared across many concurrent sessions.
  • $\bullet$
    Tail delays are disproportionately impacted by scenarios requiring large prompt recomputation or reallocation of memory-bound resources.
2.2 KV Caching as a Cross-Cutting Optimization
@@
In all popular transformer architectures, these caches are stored and used only within the lifecycle of an individual query and discarded when the query ends. Yet, KV caching can be extended to support two critical performance optimizations in production-scale serving:
@@
    Context caching (i.e.,cross-query KV cache reuse) persists KV cache segments from one query and reusing them for subsequent queries that share a common prefix. Examples include document analysis where the same document (chunk) remains constant across multiple queries, and multi-turn dialogues with a fixed system prompt or long preamble. Prefix caching reduces redundant computation during the prefill phase, directly lowering TTFT and GPU-hours per query Liu et al. (2024b); Gao et al. (2024); Jin et al. (2025b); Ren et al. (2025); Qin et al. (2025a); Jin et al. (2024); Chen et al. (2024; 2025).
@@
    Prefill–decode (PD) disaggregation (i.e.,cross-engine KV cache transfer) splits inference into a prefill stage (processing the entire input prompt) and a decode stage (autoregrssive token generation) across different GPUs or nodes. This approach reduces tail latency by maximizing the decoding speed without being interrupted by the prefill phase Zhong et al. (2024); Patel et al. (2024); Shi et al. (2025).
Roles of KV caching:  Both optimizations critically depend on the ability to move KV cache segments efficiently to and from GPU memory, CPU DRAM, NVMe storage, or over the network.
  • $\bullet$
    Context reusing stores the KV cache of a context for as long as the context will be reused, so it requires a hierarchical storage of KV caches and efficient sharing of KV caches across multiple serving engine instances.
  • $\bullet$
    PD disaggregation requires fast KV cache transfer between GPUs, often across PCIe, NVLink, or RDMA, in order to minimize the end-to-end delay of each query, which runs across the prefiller nodes and the decoder nodes.
@@ 2.3 Challenges of Efficient KV Caching
Table 1: Transfer message size vs achieved transfer throughput using RCCL transfer library UCCL Team (2025).
@@ 2.3.1 Challenge #1: I/O inefficiency under paged memory
Recent high-throughput inference engines, such as vLLM Kwon et al. (2023b) and SGLang Zheng et al. (2024), make KV cache storage and transfer even more challenging. They employ paged attention memory, dividing the attention buffer into small, fixed-size pages (typically 16–64 KB). For instance, vLLM uses 62.5-KB page in Llama-3.1-8B-Instruct model. The paged memory architecture is widely used because it improves batching and memory utilization.
However, because the pages of a KV cache are not always contiguous, the paged memory architecture dramatically increases the number of small-sized I/O operations required to persist or transfer a KV cache. Transferring such small chunks of data is known to suffer from network bandwidth underutilization and reduce throughput Kwon et al. (2025); NVIDIA Developer Forums (2020); Meta Engineering (2024). Prior work (Table 1) has shown that, on a setup with two AMD GPU nodes connected by eight Broadcom Thor-2 400Gbps NICs, the transfer size must reach at least 16 MB to saturate the available network bandwidth Zhou et al. (2025). Furthermore, prior work has shown that only transferring a data size of megabyte range (e.g., 1–2MB) can achieve 75–80% of the theoretical PCIe 5.0 bandwidth Xie et al. (2025).
@@ 2.3.2 Challenge #2: Compatible with fast-evolving inference engines
With the widespread use of AI, new LLMs and hardware accelerators are introduced at a rapid pace. In 2025, one prominent LLM was released on average every 4 days bes (2025). In response, inference engines must evolve just as quickly.
@@ 2.4 Related Work and Existing Solutions
Inference frameworks:  Since the release of vLLM Production Stack vLLM project (2025) in January 2025, there have been several open-source distributed inference stacks, including Nvidia’s Dynamo NVIDIA Corporation (2025), AIBrix Team et al. (2025), llm-d llm-d Project (2025), SGLang OME Team (2025), and KServe Contributors (2025b). They focus on easy deployment of inference engine solutions over Kubernetes, and technically, they all support various query routers based on load or prefix cache awareness and support KV caching, where LMCache is used in vLLM production stack, Dynamo, llm-d, and KServe.
@@
KV cache storage layers:  Mooncake Qin et al. (2025b), Redis Redis (2025), InfiniStore ByteDance (2025), and 3FS Contributors (2025a) provide distributed object storage or caching, but they lack an efficient “glue” layer between the inference engines to efficiently move small tensors frequently across different storage tiers, or are tightly coupled with a specific inference framework.
@@
Source code for research:  Several research proposals have open-sourced prototypes for their KV cache optimizations, including prefix caching Kwon et al. (2023a); Zheng et al. (2024); Yu et al. (2025); Gim et al. (2024); Ye et al. (2024); Lee et al. (2024); Zhao et al. (2024); Jin et al. (2024); Gao et al. (2024); Chen et al. (2025); Jin et al. (2025a); Yang et al. (2025), PD disaggregation Zhong et al. (2024); Patel et al. (2024); Shi et al. (2025), and KV cache compression Liu et al. (2024c); Jegou et al. (2024); Xiao et al. (2024a; b); Li et al. (2024); Qin et al. (2025c); Tang et al. (2024); Ge et al. (2024); Li et al. (2025); Du et al. (2025); Zhang et al. (2025). However, these prototypes are typically built on research-oriented inference framworks, such as HuggingFace Transformers, not fully enterprise-ready, or are not designed to evolve alongside the rapidly changing inference engine ecosystem, such as SGLang and vLLM.
@@ 3 Overview of LMCache
3.1 Architecture
At a high level, LMCache connects to inference engines to move KV cache between GPU memory and the storage backend. Figure 3 shows the end-to-end system. Below, we walk through two example workflows: storing and retrieving KV cache.
@@ 3.1 Architecture
3.2 Key Components
LMCache is composed of a high-performance data plane and a flexible control plane.
LMCache Worker (Data Plane).:  Each inference engine is paired with a LMCache worker. The LMCache worker module is responsible for moving KV caches between GPU memory and other tiers or other workers. It implements both KV offloading (to CPU or disk) and PD disaggregation (GPU–GPU transfer). To maximize throughput, workers employ kernel-optimized GPU buffers, asynchronous chunked I/O, and layer-wise pipelining, as elaborated in the next section. These techniques allow LMCache to sustain near-GPU-resident bandwidth even when handling small paged memory objects (16–64KB), a common configuration in vLLM and SGLang.
LMCache Controller (Control Plane).:  The controller exposes a programmable API to system operators and higher-level schedulers. It provides cache lifecycle management primitives, such as pinning KV segments, compressing and decompressing cache tensors, migrating caches between devices, and evicting low-priority entries. The controller also maintains a virtualized namespace, enabling uniform addressing of caches across heterogeneous devices.
Together, the LMCache worker and controller form a modular system that integrates seamlessly into enterprise-scale serving pipelines. The worker ensures fast data movement, while the controller provides higher-level control semantics necessary for efficient resource utilization.
@@ 4 Performance Optimizations
    modern LLM inference engines manage KV cache at the granularity of pages [^2] Each page is 16 tokens for a single layer in vLLM. , which are typically 20 KB–63 KB for popular models including Llama, Qwen, GPT-OSS etc. Such small units are inefficient for transferring, as they cannot saturate bandwidth Xie et al. (2025); Zhou et al. (2025);
@@ 4.1 Batched Operations
Configurable Chunk Size:  Rather than transferring KV cache at the page level, LMCache groups multiple pages from multiple layers into larger chunks, with a default size of 256 tokens per chunk [^3] The chunk size is configurable to different I/O speeds. . This is achieved using an intermediate GPU buffer. For storing, multiple pages (16 by default) are first copied from GPU memory into the buffer, then offloaded collectively to lower-tier storage (e.g., CPU memory) and saved at the granularity of chunks rather than individual pages. For loading, chunks are first retrieved from the storage layer into the GPU buffer and subsequently split into pages before being placed in the serving engine’s internal GPU memory. LMCache implements a customized CUDA kernel to accelerate the memory copy.
Batched Store/load Operations:  LMCache supports storing and loading KV cache across multiple storage devices, including local CPU memory, local disks, and remote disks. In practice, KV caches often need to be transferred to different devices in parallel – for example, storing the KV cache of a hot context in CPU memory while simultaneously offloading a cold context to local disk. A naive approach that stores the two KV caches sequentially underutilizes available bandwidth, since transferring to CPU memory leaves the disk bandwidth idle. To address this, LMCache batches storage and loading queries across different tiers, thereby aggregating operations and fully utilizing the bandwidth between the GPU and multiple storage tiers.
Delayed Decode KV Cache Storing:  LMCache also supports saving newly generated KV cache during decoding. In a naive design, each newly generated page is immediately offloaded to lower-tier storage, incurring frequent small writes. Instead, LMCache aggregates multiple pages and stores them together as a chunk, improving I/O efficiency.

### v2

@@
LMCache: An Efficient KV Cache Layer for Enterprise-Scale LLM Inference
Yuhan Liu12* Jiayi Yao12* Yihua Cheng1* Yuwei An1 Xiaokun Chen1 Shaoting Feng12
Yuyang Huang12 Samuel Shen1 Rui Zhang1 Kuntai Du1 Junchen Jiang1
1Tensormesh Inc. 2University of Chicago
footnotetext: *Equal contribution
@@ 1 Introduction
Today, large-language model (LLM) inference has outpaced training in growth. LLM inference powers millions of applications, from interactive customer support and code generation to retrieval-based document analysis and agentic workflows. To build the systems for LLM inference, KV cache, the intermediate states of large language model (LLM) inference, has now become a de facto optimization to make inference faster.
Traditionally, KV cache has been used during new token generation to skip re-computing the KV cache of the input prompts, within a single query. Thus, every query is processed independently by one instance of the inference engine, and a LLM query’s lifecycle, including computation and I/O operations, happens in the GPUs and the GPU memory of one inference engine.
Figure 1: Upper: LMCache is being used by more and more users over time. Middle: LMCache is being used to store KV cache of larger size. Bottom: LMCache’s docker image pull counts have continued to increase.
However, recent trends has advanced to propose moving KV cache outside the GPU memory, including two emerging directions:
Cross-query caching to avoid redundant compute:  KV cache can be persisted (e.g., to lower-tier storage devices) beyond the lifecycle of a query to avoid any re-computation of the shared prefix for another query.
Prefill–decode disaggregation for higher utilization:  There is an arising trend to decouple the prefill and decode phase on different GPUs, to ensure that latency-sensitive decoding phase is not affected by throughput-oriented prefill phase. Such prefill-decode (PD) diaggregation requires KV cache produced by the prefill GPUs to be transferred to the decode GPUs.
Our real-world usage statistics confirm the trend of moving KV cache out from the GPU memory. As will be discussed in detail soon (§2.2), over time, the total size of KV cache that is stored by the users continues to grow, and it far exceeds the capacity of the GPU memory. This suggests that KV cache may need to be frequently evicted from GPU memory as more and more requests come in. Also, in order for it to be reused by another query, it needs to be offloaded from GPU memory, and then loaded back to GPU memory for reuse.
@@
To make these ideas practical, the LLM inference systems must be augmented with new KV cache semantics. In particular, inference engines should support the new interface that extracts KV caches from a normal inference call and re-loads KV caches into subsequent queries on demand. The system also must allow the extracted KV caches to be stored persistently and transferred across distributed inference engines. Most importantly, for such interface extensions to be practical, KV cache extraction, re-loading, storage, and transfer must be efficient, and the new interfaces must remain compatible with rapidly evolving inference engines such as vLLM vllm and SGLang zheng2024sglangefficientexecutionstructured.
@@
Beyond quantitative gains, LMCache has seen adoption across several enterprises and open-source projects, providing useful insights and lessons in KV cache-driver optimizations at production scale. This includes surprising gain in latency brought by remote storage backend, reduction in prefix cache hit rate because of context truncation, and that the flexibility to evolve faster and integration smoothness is more important than the language performance itself.
The remainder of this paper details the motivation (§2) and challenges (§3), LMCache architecture and key design choices (§4, §5, and §6), deployment experiences (§9) and experimental evaluation (§8).
2 Motivation and Real-world Usage Statistics
2.1 KV Cache in LLM Inference
@@
Nowadays, the contexts have grow longer and longer, and people have started to augment inference with background knowledge. Given this trend, it is popular to share KV cache across different user queries to reduce the redundant computations for the long contexts or background knowledge.
Figure 3: Weekly growth of KV cache size, including portions that fit in GPU memory and those that exceed it.
2.2 Real-world Usage Statistics
KV Cache Size Exceeds GPU memory:  Although KV cache has been kept inside GPU memory for all traditional LLM inference systems, we observe that the required size of KV cache is now far exceeding the GPU memory capacity, from our real-world usage statistics, gathered by usage tracker voluntarily turned on by users.
Figure 3 shows the weekly growth of KV cache size over the past five weeks, for caches that fit within GPU memory (green) and those that exceed GPU memory capacity (blue). The portion of KV cache that no longer fits in GPU memory has increased significantly over time, showing that GPU memory alone is insufficient for storing all caches. To enable KV cache reuse across queries, especially those generated long before reuse, it becomes necessary to move KV cache out of GPU memory, for instance, by offloading it to CPU memory or other storage tiers.
Reuse per Token has Greatly Increased:  We also observe that reuses per token has greatly increased over time.
As shown on the left side of Figure 4, where we plot the ratio between reused tokens and all stored tokens, beyond the GPU memory, plotted with top-10 users. We denote this ratio as reuses per token. The reuses per token has grown significantly over the past several weeks, which indicates that tokens that cannot fit inside GPU memory are being more and more frequently reused by inference. This suggests that more and more tokens need to be loaded back to GPU memory.
On the right hand side, the figure shows the distribution of reuse per token for different users over the past week. More than 19% of users reuse stored tokens for more than 1.5 times, suggesting the trend of users accessing a token multiple times after it is stored.
Figure 4: Left: Average reuse per token for top users. Right: Distributions of average reuse per token across different users.
2.3 Need an Efficient KV Caching Layer for Moving KV Cache
From the above observations from the statistics gathered in real-world deployments, we find two important trends of KV caching. First, the KV cache that cannot simply fits in GPU memory keeps growing, potentially due to growing length of the contexts or larger amount of user traffic. Second, reuses per token stored beyond the GPU memory has also increased over time. Both trends suggest that we need to move KV cache out of GPU memory. Specifically, in the current industry, two scenarios which move KV cache out from GPU exist:
@@ 2.3 Need an Efficient KV Caching Layer for Moving KV Cache
    Context caching (i.e.,cross-query KV cache reuse) persists KV cache segments from one query and reusing them for subsequent queries that share a common prefix. Examples include document analysis where the same document (chunk) remains constant across multiple queries, and multi-turn dialogues with a fixed system prompt or long preamble. Prefix caching reduces redundant computation during the prefill phase, directly lowering TTFT and GPU-hours per query liu2024cachegenkvcachecompression; gao2024costefficientlargelanguagemodel; cake; 10.1145/3719330.3721230; mooncake; jin2024ragcacheefficientknowledgecaching; chen2024kvdirectdistributeddisaggregatedllm; impress.
@@
    Prefill–decode (PD) disaggregation (i.e.,cross-engine KV cache transfer) splits inference into a prefill stage (processing the entire input prompt) and a decode stage (autoregrssive token generation) across different GPUs or nodes. This approach reduces tail latency by maximizing the decoding speed without being interrupted by the prefill phase distserve; splitwise; shi2025nexusproactiveintragpudisaggregationprefill.
However, there lacks a library to support efficient extraction and loading from and to the GPU memory due to the system challenges as discussed soon (3).
3 Challenges of Efficient KV Caching and Related Work
@@ 3.1 Challenges of Efficient KV Caching
Table 1: Transfer message size vs achieved transfer throughput using RCCL transfer library UCCL2025KVTransferEngine.
@@ 3.1.1 Challenge #1: I/O inefficiency under paged memory
Recent high-throughput inference engines, such as vLLM vllm and SGLang zheng2024sglangefficientexecutionstructured, make KV cache storage and transfer even more challenging. They employ paged attention memory, dividing the attention buffer into small, fixed-size pages (typically 16–64 KB). For instance, vLLM uses 62.5-KB page in Llama-3.1-8B-Instruct model. The paged memory architecture is widely used because it improves batching and memory utilization.
However, because the pages of a KV cache are not always contiguous, the paged memory architecture dramatically increases the number of small-sized I/O operations required to persist or transfer a KV cache. Transferring such small chunks of data is known to suffer from network bandwidth underutilization and reduce throughput kwon2023demystifyingnccl; nvidia2020smalltransfer; facebook2024roce. Prior work (Table 1) has shown that, on a setup with two AMD GPU nodes connected by eight Broadcom Thor-2 400Gbps NICs, the transfer size must reach at least 16 MB to saturate the available network bandwidth uccl_transport. Furthermore, prior work has shown that only transferring a data size of megabyte range (e.g., 1–2MB) can achieve 75–80% of the theoretical PCIe 5.0 bandwidth strata.
@@ 3.1.2 Challenge #2: Compatible with fast-evolving inference engines
With the widespread use of AI, new LLMs and hardware accelerators are introduced at a rapid pace. In 2025, one prominent LLM was released on average every 4 days best44llms2025. In response, inference engines must evolve just as quickly.
@@ 3.2 Related Work and Existing Solutions
Inference frameworks:  Since the release of vLLM Production Stack vllm-production-stack2025 in January 2025, there have been several open-source distributed inference stacks, including Nvidia’s Dynamo aiDynamoDynamo2025, AIBrix theaibrixteam2025aibrixscalablecosteffectivelarge, llm-d llmDllmD2025, SGLang OME ome, and KServe KServe2025. They focus on easy deployment of inference engine solutions over Kubernetes, and technically, they all support various query routers based on load or prefix cache awareness and support KV caching, where LMCache is used in vLLM production stack, Dynamo, llm-d, and KServe.
@@
KV cache storage layers:  Mooncake qin2025mooncakekvcachecentricdisaggregatedarchitecture, Redis redis-enterprise-software-references, InfiniStore bytedanceInfiniStore2025, and 3FS DeepSeek2025ThreeFS provide distributed object storage or caching, but they lack an efficient “glue” layer between the inference engines to efficiently move small tensors frequently across different storage tiers, or are tightly coupled with a specific inference framework.
@@
Source code for research:  Several research proposals have open-sourced prototypes for their KV cache optimizations, including prefix caching vllm; zheng2024sglangefficientexecutionstructured; pensieve; prompt-cache; chunkattention; infinigen; blendserve; jin2024ragcacheefficientknowledgecaching; gao2024costefficientlargelanguagemodel; impress; cake; kvshare, PD disaggregation distserve; splitwise; shi2025nexusproactiveintragpudisaggregationprefill, and KV cache compression liu2024kivi; kvpress2024; xiao2024duoattentionefficientlongcontextllm; xiao2024efficientstreaminglanguagemodels; snapkv; qin2025cakecascadingadaptivekv; quest; ge2024modeltellsdiscardadaptive; commvq; bitdecoding; pqcache. However, these prototypes are typically built on research-oriented inference framworks, such as HuggingFace Transformers, not fully enterprise-ready, or are not designed to evolve alongside the rapidly changing inference engine ecosystem, such as SGLang and vLLM.
@@ 4 Overview of LMCache
Figure 6 shows the end-to-end system. Below, we walk through two example workflows: storing and retrieving KV cache.
@@ 5 Performance Optimizations
    Modern LLM inference engines manage KV cache at the granularity of pages [^2] Each page is 16 tokens for a single layer in vLLM. , which are typically 20 KB–63 KB for popular models including Llama, Qwen, GPT-OSS etc. Such small units are inefficient for transferring, as they cannot saturate bandwidth strata; uccl_transport.
@@ 5.1 Batched Operations
Configurable Chunk Size:  Rather than transferring KV cache at the page level, LMCache groups multiple pages from multiple layers into larger chunks, with a default size of 256 tokens per chunk [^3] The chunk size is configurable to different I/O speeds. . This is achieved using an intermediate streaming GPU buffer. For storing, the KV cache are first copied from the scattered paged GPU memory into a contiguous streaming buffer with a customized CUDA kernel, then offloaded collectively to lower-tier storage (e.g., CPU memory) with DMA engines at the granularity of chunks rather than individual pages. For loading, chunks are first retrieved from the storage layer into the GPU buffer with DMA engines and subsequently split into paged memory with CUDA kernels.
Parallel Store/load Operations:  LMCache supports parallel storage and retrieval of KV caches across multiple storage tiers, including local CPU DRAM or disks, remote CPU DRAM or disks, and object storage (e.g., S3). In practical LLM serving workloads, KV caches often need to be migrated across devices concurrently—for instance, transferring the KV cache of a hot context from GPU to CPU memory while simultaneously offloading a cold context from CPU memory to local disk. To maximize link utilization, LMCache ’s store and load APIs accept multiple source and destination devices, enabling concurrent data movement across heterogeneous links. Moreover, these operations can be executed in parallel when the interconnect supports full-duplex communication (e.g., PCIe).
Delayed Decode KV Cache Storing:  LMCache also supports storing newly generated KV caches during decoding. Instead of offloading each token’s KV cache immediately—a naive approach that triggers frequent small writes—LMCache buffers KV caches and performs batched storage once a predefined number of tokens (i.e., a chunk) have been generated. This chunk-based delayed storing strategy reduces write frequency, minimizes I/O overhead, and significantly improves overall storage throughput.